# Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young M4 + [T7 + T8] hierarchical triple

### v1

@@ 4.1 M dwarf companions
The Gaia Renormalized Unit Weight Error (RUWE) values for L 122-88 A and UPM J10403551 A (1.32 and 1.47, respectively) suggest these objects may potentially be unresolved binaries. However, their optical spectra appear consistent with normal M dwarfs (Fig. 3), indicating that if they are indeed binaries, each component likely possesses a similar spectral classification. Under such circumstances, given their identical distances, metallicity constraints derived from model fitting, whether applied to their composite spectra or to individual components, would yield consistent results. In this section, we examine their spectra as single objects.
@@ 4.1.1 L 122-88 A
We compared the optical spectra (Fig. 3) of L 122-88 A to those M0-M9 dwarf templates from Bochanski et al. (2007), and found it closely matches an M2 dwarf template. However, it exhibits slightly weaker TiO absorption-bands around 710 nm, which indicates a mildly metal-poor composition (e.g., Lépine et al., 2007; Jao et al., 2008; Zhang, 2019).
Analysis of Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) data using the TESSILATOR software (Binks & Günther, 2024) reveals a rotation signal of 2.4750.013 days in sector 2 data. Applying gyrochronology to this rotation period using GPgyro (Lu et al., 2024) yields an age estimate of 0.0290.010 Gyr. However, there is no evidence for this signal in other TESS sectors 28, 29, 68, and 69. Examination of the TESS data sector 2 release notes222https://ntrs.nasa.gov/citations/20190001284 for TIC 271503151 (L 122-88 A) does not suggest any particular issue with the sector 2 observation but for a V=13 object we would expect the signal to be visible in other sectors for a reliable detection. Marocco et al. (2024) suggested a significantly older kinematic age of 2.5-9.9 Gyr for L 122-88 A.
@@ 4.1.2 UPM J10403551 A
We compared the optical spectrum (Fig. 3) of UPM J10403551 A to those M0-M9 dwarf templates from Bochanski et al. (2007), and found it closely matches that of an M4 dwarf template. The presence of H emission in its spectrum indicates that it is an active M4 dwarf, which is also suggestive of a relatively young age. We measured the H equivalent width (H EW) of UPM J10403551 A to be between 2.33 and 3.78 Å. Applying the H activity-age correlation for M3-6 dwarfs from Kiman et al. (2021) (their fig. 6), this H EW range corresponds to an estimated age of 0.3-2.0 Gyr. The TESS data for UPM J10403551 A shows no indication of activity or any periodic signal.
@@ 4.1.3 Atmospheric properties
The optical spectrum is more sensitive to of M (sub)dwarfs, while the NIR spectrum is more sensitive to [Fe/H]. Therefore, the full optical-NIR spectral profile of M subdwarfs can have better constraints of their and [Fe/H]. To derive the atmospheric parameters of L 122-88 A, we fitted its X-shooter spectrum (300-2500 nm) to BT-Settl models with 2600 K K, [Fe/H] 0.5, and log = 5.0 (Allard, 2014). We used a typical log = 5.0 for M2-M4 dwarfs as log has relatively small affection on their spectra compared to and [Fe/H]. We applied smoothing to the X-shooter spectrum to enhance the SNR, though this degraded the resolving power to 1000. The BT-Settl model spectra were similarly degraded to match the resolution of the processed X-shooter spectrum.
@@ 4.2.1 L 122-88 B
HIP 73786 B exhibits a weak flux excess on the left-hand side of the -band peak (980-1080 nm), which is possibly caused by subsolar metallicity. This feature is also visible in spectra of other T subdwarfs with [Fe/H] and becomes more significant at lower metallicity (e.g., Pinfield et al., 2012; Burningham et al., 2014). The absence of the -band flux excess feature in L 122-88 B indicates a metallicity higher than that of T subdwarfs, which usually have [Fe/H] (e.g., fig. 9, Zhang et al., 2017b). This is still consistent with the metallicity derived from L 122-88 A.
To further investigate the mildly metal-poor nature of L 122-88 B, we compared its absolute magnitudes in the and -bands, computed using the Gaia parallax of the primary, to those of known T dwarfs and subdwarfs (Fig. 7). L 122-88 B exhibits slightly fainter and compared to typical T5 dwarfs, while its remains similar. This pattern is consistent with the behaviour of T5.5-9 subdwarfs, which show fainter and -band absolute magnitudes than their solar-metallicity counterparts, but indistinguishable -band absolute magnitudes (Zhang et al., 2019). Without considering its mildly metal-poor nature, L 122-88 B might be misclassified as a T6.51 dwarf based on spectral type versus absolute magnitude correlations (Fig. 7). Indeed, Marocco et al. (2024) classified it as a T6 dwarf based on spectral type versus colour correlations from Kirkpatrick et al. (2021).
We also estimated the distance of L 122-88 B by the spectral type – absolute magnitude correlations (fig. 10, Zhang et al., 2019). The average distance estimated from its magnitudes is 45 pc, which is consistent to the Gaia distance of L 122-88 A (33.1060.014) within the uncertainty. Of course, L 122-88 B exhibits a over estimated spectral distance due to its mildly metal-poor composition. This metallicity deficiency causes its actual absolute magnitudes to be fainter than the average values for T5 dwarfs used in the calculation.
@@ 4.2.2 UPM J10403551 B: a probable unresolved binary
The NIR spectrum of UPM J10403551 B closely resembles that of the T7 dwarf standard 2MASS J0727182+171001 (J0727+17; Burgasser et al., 2002, 2006). However, Fig. 4 reveals flux excesses in the and bands relative to the T7 standard when normalized at the -band peak. When rescaling the spectrum of UPM J1040-3551 B to better align with the T7 standard across , , and bands, we observe a 13% flux suppression in the band, a phenomenon neither previously observed nor explained in any single BD.
@@
Several features of evidence suggest that UPM J10403551 B is not a T subdwarf: (1) it exhibits -band flux excess rather than the suppression typical of late-type T subdwarfs, (2) it has a relatively young age of 0.3-2.0 Gyr inferred from its primary (Section 4.1.2), and (3) it has solar metallicity as indicated by its primary’s spectrum (Fig. 3). While the -band flux excess could be attributed to its young age (e.g., Burningham et al., 2011), this explanation fails to account for the observed and band flux excesses.
@@
Unresolved binary BDs may show unusual spectral features and colours, and flux excesses if have been treated as single objects (Burgasser et al., 2010; Zhang et al., 2010; Marocco et al., 2015; Zhang et al., 2018b; Kirkpatrick et al., 2021). Figure 7 demonstrates that UPM J10403551 B has significantly brighter and band absolute magnitudes than typical T7 dwarfs. Such over-luminosity in field ultracool dwarfs often indicates the presence of an unresolved companion.
@@
To test this hypothesis, we conducted a binary spectral fitting analysis using single spectral templates in the Spex Prism Library Analysis Toolkit (Burgasser & Splat Development Team, 2017, SPLAT;). We selected 131 T4–T9 dwarfs with low-resolution ( 150) NIR spectra previously obtained with the SpeX spectrograph (Rayner et al., 2003) on the 3m NASA Infrared Telescope Facility. These templates exclude known or suspected binaries and spectra with very low S/N 15. Spectra were interpolated and resampled onto a common wavelength scale spanning 0.95–2.5 m at a constant resolution = 200, classified by comparison to NIR spectral standards (Burgasser et al., 2006), and then flux-calibrated using the Kirkpatrick et al. (2021) absolute MKO -band magnitude/spectral type relation. We then co-added pairs of spectra to create 6,715 unique binary templates. We compared both single and binary templates to the equivalently interpolated SOAR spectrum of J10403551 B following procedures described in (Burgasser et al., 2010), masking out regions of strong telluric absorption between 1.35–1.42 m and 1.80–1.95 m. Figure 8 compares the best fit single template from this sample, the T7 dwarf WISE J004024.88+090054.8 (Mace et al., 2013, J0040+0900;) to the best-fit binary template composed of J0040+0900 and the T8 dwarf ULAS J102940.52+093514.6 (Burningham et al., 2013; Thompson et al., 2013, J1029+0935;). The latter is a superior fit, notably providing a better reproduction of the , , and -band peaks. The difference in reduced between the single fit (10.6) and the binary fit (7.6) implies a 2% likelihood of equivalence based on the F-test statistic. However, we caution that other physical parameters, such as youth/low surface gravity, could also explain the modest deviations in the single template fit and may not be fully accounted for in the existing SPLAT template sample. The spectral fitting therefore provides supporting but not conclusive evidence of unresolved multiplicity.
@@
To further investigate this binary hypothesis, we decomposed the absolute magnitudes of UPM J10403551 B into T7 and T8 components. Using the spectral type versus absolute magnitude correlations for L and T dwarfs (Fig. 7), we derived , , and differences between T7 and T8 dwarfs of 1.258, 1.248, 0.907, and 0.462 mag, respectively. The decomposed , , and values demonstrate improved alignment with these correlations, though they remain slightly brighter than average—contrary to the dimmer characteristics of late T subdwarfs. UPM J10403551 B appears underluminous as an unresolved binary only in the -band, which may be attributed to its young age. For instance, the young T5.5 dwarf SDSSp J111010.01+011613.1 (Geballe et al., 2002) is relatively faint in the -band, with a colour 0.34 mag bluer than the average for field T5.5 dwarfs (fig. 8i of Primeval VI, ).
@@
The independent spectroscopic distance of UPM J10403551 B as a single T7 dwarf is 20 pc, which is 21% shorter than the Gaia distance of the M4 primary (25.2830.013 pc). We also decomposed the band apparent magnitudes of UPM J10403551 B as T7 + T8 close binary. The average spectroscopic distance of UPM J10403551 Ba as a T7 dwarf is 21 pc, which is 17% smaller than its UPM J10403551 A but consistent within the uncertainty. The spectral distance of UPM J10403551 B is under estimated due to its young age, which allows it to have brighter absolute magnitudes than the average values for T7 dwarfs used in the calculation.
@@
Both the spectral fitting and the decomposition of absolute magnitudes strongly suggest that UPM J10403551 B is likely an unresolved T7 + T8 binary (UPM J10403551 Ba & Bb). UPM J10403551 ABab might be the third known triple system comprising a close binary of T dwarfs with a wide stellar companion, similar to Indi ABab (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010) or Gliese 229 ABab (Xuan et al., 2024).
@@
UPM J10403551 B is detected but not resolved in the Visible and Infrared Survey Telescope for Astronomy (VISTA) Hemisphere Survey (VHS; McMahon et al., 2013) and the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019b). The VHS and DECaLS images have pixel sizes of 034 and 026, respectively. This allows us to place upper limits on the orbital semi-major axis ( au) and period ( yr) of the system. At a distance of 25 pc, the components can potentially can be resolved at 1.15 m by the Near-Infrared Camera (NIRCam) of the James Webb Space Telescope (JWST), provided the orbital period exceeds six years.
@@ 4.3 Binary properties
To estimate the masses of the T dwarf companions, we employed the evolutionary tracks of from the Sonora-Bobcat models (Marley et al., 2021), using their estimated ages and the average of T dwarfs with corresponding spectral types. As illustrated in Fig. 9, L 122-88 B has an estimated mass of 0.05–0.067 M for an age of 2.5-9.9 Gyr, while UPM J10403551 Ba and Bb have estimated masses of 0.012–0.04 M and 0.007–0.02 M⊙, respectively, for an age of 0.3-2.0 Gyr. Using the derived distances, projected separations, and masses of all components, we calculated the binding energies of L 122-88 AB and UPM J10403551 ABab to be J and J, respectively.
@@
We found that L 122-88 AB and UPM J10403551 ABab have values of 1.02 pc and 0.87 pc, respectively. Their projected separations correspond to 0.034 and 0.009 , respectively, which are far smaller than their respective values. This analysis strongly suggests that both systems are gravitationally stable. The spectroscopic distances of L 122-88 B and UPM J10403551 B are also consistent to Gaia parallax distances of their primaries (see Table 1).

### v3

@@ 4.1 M dwarf companions
The Gaia Renormalized Unit Weight Error (RUWE) values for L 122-88 A and UPM J10403551 A (1.32 and 1.47, respectively) suggest that these objects may potentially be unresolved binaries. However, their optical spectra appear consistent with normal M dwarfs (Fig. 3), indicating that if they are indeed binaries, each component likely possesses a similar spectral classification. Under such circumstances, given their identical distances, metallicity constraints derived from model fitting, whether applied to their composite spectra or to individual components, would yield consistent results. In this section, we first examine their spectra as single objects, then discuss their binarity in Section 4.1.4.
@@ 4.1.1 L 122-88 A
We compared the optical spectra (Fig. 3) of L 122-88 A to those M0-M9 dwarf templates from Bochanski et al. (2007), and found that it closely matches an M2 dwarf template. However, it exhibits slightly weaker TiO absorption bands around 710 nm, which indicates a mildly metal-poor composition (e.g., Lépine et al., 2007; Jao et al., 2008; Zhang, 2019).
Analysis of data from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) using the TESSILATOR software (Binks & Günther, 2024) reveals a rotation signal of 2.475 0.013 d in sector 2 data. Applying gyrochronology to this rotation period using GPgyro (Lu et al., 2024) yields an age estimate of 0.0290.010 Gyr. However, there is no evidence for this signal in other TESS sectors (28, 29, 68, and 69). Examination of the TESS data sector 2 release notes222https://ntrs.nasa.gov/citations/20190001284 for TIC 271503151 (L 122-88 A) does not suggest any particular issue with the sector 2 observation but for a = 13 object we would expect the signal to be visible in other sectors for a reliable detection. Marocco et al. (2024) suggested a significantly older kinematic age of 2.5-9.9 Gyr for L 122-88 A.
@@ 4.1.2 UPM J10403551 A
We compared the optical spectrum (Fig. 3) of UPM J10403551 A to those M0–M9 dwarf templates from Bochanski et al. (2007), and found it closely matches that of an M4 dwarf template. The presence of H emission in its spectrum indicates that it is an active M4 dwarf, which is also suggestive of a relatively young age. We measured the H equivalent width (H EW) of UPM J10403551 A to be between 2.33 and 3.78 Å. Applying the H activity–age correlation for M3–M6 dwarfs from Kiman et al. (2021) (their fig. 6), this H EW range corresponds to an estimated age of 0.3-2.0 Gyr. The TESS data for UPM J10403551 A show no indication of activity or any periodic signal.
@@ 4.1.3 Atmospheric properties
The optical spectrum is more sensitive to of M (sub)dwarfs, while the NIR spectrum is more sensitive to [Fe/H]. Therefore, the full optical–NIR spectral profile of M subdwarfs can have better constraints of their and [Fe/H]. To derive the atmospheric parameters of L 122-88 A, we fitted its X-shooter spectrum (300–2500 nm) to BT-Settl models with 2600 K K, [Fe/H] 0.5, and log = 5.0 (Allard, 2014). We used a typical log = 5.0 for M2–M4 dwarfs as log has relatively small effect on their spectra compared to and [Fe/H]. We applied smoothing to the X-shooter spectrum to enhance the SNR, though this degraded the resolving power to 1000. The BT-Settl model spectra were similarly degraded to match the resolution of the processed X-shooter spectrum.
@@
4.1.4 UPM J10403551 A: a likely near equal-mass binary
Single stars are expected to have RUWE 1.0. Therefore, RUWE 1.4 is usually used as an indicator of non-single object. Nearby M dwarfs and earlier type stars predominantly exhibit RUWE values of 1.2 0.3 and 1.0 0.2, respectively (Sozzetti, 2023). The RUWE = 1.47 of UPM J10403551 A suggests it is likely an unresolved binary. For unresolved equal-mass binaries, the combined magnitude appears 0.75 mag brighter than each individual component, providing a method to test the binarity hypothesis.
UPM J10403551 A exhibits a typical (3200 50 K) for M4 dwarfs (see fig. 4 of Zhang et al., 2018a). Notably, model spectral fitting would yield identical temperature estimates regardless of whether the object is a single star or an equal-mass binary. According to evolutionary models (Baraffe et al., 2015), a solar-metallicity star with K and age of 0.5-5 Gyr should have a mass of 0.174 M⊙, radius of 0.195 R⊙, and absolute magnitudes of , , and . The observed absolute magnitudes of UPM J10403551 A (, , and ) are systematically brighter by 0.69, 0.73, and 0.70 mag, respectively, than predicted by models for its temperature (see Fig. 7). This consistent 0.7 mag excess strongly supports the classification of UPM J10403551 A as a near equal-mass binary system (UPM J10403551 Aab), with components Aa and Ab differing by only 0.1 mag ( 25 K) and both can be classified as M4 dwarfs.
As a single star, UPM J10403551 A would have a mass of M⊙ according to the mass- correlation (Mann et al., 2019). A solar-metallicity star with this mass and age of 0.5-5 Gyr should have a K, significantly higher than the model spectral fitting result. With of 3200 K and 3175 K, UPM J10403551 Aa & Ab should have a mass around 0.174 0.022 M⊙ and 0.165 0.020 M⊙, respectively, according to evolutionary models (Baraffe et al., 2015).
For comparison, we also evaluated L 122-88 A by comparing its observed , , and values to evolutionary model predictions for a solar-metallicity star with K and age of 8 Gyr (see Fig. 7). The resulting magnitude differences of approximately 0.11 mag are insufficient to suggest binarity in this case.
@@ 4.2.1 L 122-88 B
HIP 73786 B exhibits a weak flux excess on the left-hand side of the -band peak (980-1080 nm), which is possibly caused by subsolar metallicity. This feature is also visible in spectra of other T subdwarfs with [Fe/H] and becomes more significant at lower metallicity (e.g., Pinfield et al., 2012; Burningham et al., 2014). The absence of the -band flux excess feature in L 122-88 B indicates a metallicity higher than that of T subdwarfs, which usually have [Fe/H] (e.g., fig. 9 of Zhang et al., 2017b). This is still consistent with the metallicity derived from L 122-88 A.
To further investigate the mildly metal-poor nature of L 122-88 B, we compared its absolute magnitudes in the and bands, computed using the Gaia parallax of the primary, to those of known T dwarfs and subdwarfs (Fig. 8). L 122-88 B exhibits slightly fainter and compared to typical T5 dwarfs, while its remains similar. This pattern is consistent with the behaviour of T5.5-9 subdwarfs, which show fainter and -band absolute magnitudes than their solar-metallicity counterparts, but indistinguishable -band absolute magnitudes (Zhang et al., 2019). Without considering its mildly metal-poor nature, L 122-88 B might be misclassified as a T6.5 1 dwarf based on spectral type versus absolute magnitude correlations (Fig. 8). Indeed, Marocco et al. (2024) classified it as a T6 dwarf based on spectral type versus colour correlations from Kirkpatrick et al. (2021).
We also estimated the distance of L 122-88 B by the spectral type – absolute magnitude correlations (fig. 10 of Zhang et al., 2019). The average distance estimated from its magnitudes is 45 pc, which is consistent to the Gaia distance of L 122-88 A (33.106 0.014) within the uncertainty. Of course, L 122-88 B exhibits an over estimated spectral distance due to its mildly metal-poor composition. This metallicity deficiency causes its actual absolute magnitudes to be fainter than the average values for T5 dwarfs used in the calculation.
@@ 4.2.2 UPM J10403551 B: a probable unresolved binary
The NIR spectrum of UPM J10403551 B closely resembles that of the T7 dwarf standard 2MASS J0727182+171001 (J0727+17; Burgasser et al., 2002, 2006). However, Fig. 4 reveals flux excesses in the and bands relative to the T7 standard when normalized at the -band peak. When rescaling the spectrum of UPM J10403551 B to better align with the T7 standard across , , and bands, we observe a 13% flux suppression in the band, a phenomenon neither previously observed nor explained in any single BD.
@@
Several features of evidence suggest that UPM J10403551 B is not a T subdwarf: (1) it exhibits -band flux excess rather than the suppression typical of late-type T subdwarfs; (2) it has a relatively young age of 0.3-2.0 Gyr inferred from its primary (Section 4.1.2); and (3) it has solar metallicity as indicated by its primary’s spectrum (Fig. 3). While the -band flux excess could be attributed to its young age (e.g., Burningham et al., 2011), this explanation fails to account for the observed - and -band flux excesses.
@@
Unresolved binary BDs may show unusual spectral features and colours, and flux excesses if they have been treated as single objects (Burgasser et al., 2010; Zhang et al., 2010; Marocco et al., 2015; Zhang et al., 2018b; Kirkpatrick et al., 2021). Figure 8 demonstrates that UPM J10403551 B has significantly brighter -, -, and -band absolute magnitudes than typical T7 dwarfs. Such overluminosity in field ultracool dwarfs often indicates the presence of an unresolved companion.
@@
To test this hypothesis, we conducted a binary spectral fitting analysis using single spectral templates in the Spex Prism Library Analysis Toolkit (SPLAT; Burgasser & Splat Development Team, 2017). We selected 131 T4–T9 dwarfs with low-resolution ( 150) NIR spectra previously obtained with the Spectral cross-disperser (SpeX) spectrograph (Rayner et al., 2003) on the 3m NASA Infrared Telescope Facility. These templates exclude known or suspected binaries and spectra with very low S/N 15. Spectra were interpolated and resampled on to a common wavelength scale spanning 0.95–2.5 m at a constant resolution = 200, classified by comparison to NIR spectral standards (Burgasser et al., 2006), and then flux-calibrated using the Kirkpatrick et al. (2021) absolute MKO -band magnitude/spectral type relation. We then co-added pairs of spectra to create 6,715 unique binary templates. We compared both single and binary templates to the equivalently interpolated SOAR spectrum of J10403551 B following procedures described in Burgasser et al. (2010), masking out regions of strong telluric absorption between 1.35–1.42 m and 1.80–1.95 m. Figure 9 compares the best fitting single template from this sample, the T7 dwarf WISE J004024.88+090054.8 (Mace et al., 2013, J0040+0900;) to the best-fitting binary template composed of J0040+0900 and the T8 dwarf ULAS J102940.52+093514.6 (J1029+0935; Burningham et al., 2013; Thompson et al., 2013). The latter is a superior fit, notably providing a better reproduction of the , , and -band peaks. The difference in reduced between the single fit (10.6) and the binary fit (7.6) implies a 2% likelihood of equivalence based on the F-test statistic. However, we caution that other physical parameters, such as youth/low surface gravity, could also explain the modest deviations in the single template fit and may not be fully accounted for in the existing SPLAT template sample. The spectral fitting therefore provides supporting but not conclusive evidence of unresolved multiplicity.
@@
To further investigate this binary hypothesis, we decomposed the absolute magnitudes of UPM J10403551 B into T7 and T8 components. Using the spectral type versus absolute magnitude correlations for L and T dwarfs (Fig. 8), we derived , , and differences between T7 and T8 dwarfs of 1.258, 1.248, 0.907, and 0.462 mag, respectively. The decomposed , , and values demonstrate improved alignment with these correlations, though they remain slightly brighter than average – contrary to the dimmer characteristics of late T subdwarfs. UPM J10403551 B appears underluminous as an unresolved binary only in the band, which may be attributed to its young age. For instance, the young T5.5 dwarf SDSSp J111010.01+011613.1 (Geballe et al., 2002) is relatively faint in the band, with a colour 0.34 mag bluer than the average for field T5.5 dwarfs (fig. 8i of Zhang et al., 2019).
@@
The independent spectroscopic distance of UPM J10403551 B as a single T7 dwarf is 20 pc, which is 21 percent shorter than the Gaia distance of the M4 primary (25.2830.013 pc). We also decomposed the -, -, -, -band apparent magnitudes of UPM J10403551 B as T7 + T8 close binary. The average spectroscopic distance of UPM J10403551 Ba as a T7 dwarf is 21 pc, which is 17 percent smaller than its UPM J10403551 A but consistent within the uncertainty. The spectral distance of UPM J10403551 B is under estimated due to its young age, which allows it to have brighter absolute magnitudes than the average values for T7 dwarfs used in the calculation.
@@
Both the spectral fitting and the decomposition of absolute magnitudes strongly suggest that UPM J10403551 B is likely an unresolved T7 + T8 binary (UPM J10403551 Ba & Bb). UPM J10403551 AabBab is the first known quadruple system comprising a close binary of T dwarfs in a wide orbit around a close binary of stars. It is the third identified system featuring a close binary of T dwarfs with wide stellar companion, following Indi ABab (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010) and Gliese 229 ABab (Xuan et al., 2024).
@@
UPM J10403551 B is detected but not resolved in the Visible and Infrared Survey Telescope for Astronomy (VISTA) Hemisphere Survey (VHS; McMahon et al., 2013) and the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019). The VHS and DECaLS images have pixel sizes of 0.34 and 0.26 arcsec, respectively. This allows us to place upper limits on the orbital semimajor axis ( au) and period ( yr) of the system. At a distance of 25 pc, the components can potentially be resolved at 1.15 m by the Near-Infrared Camera of the JWST, provided the orbital period exceeds six years.
@@ 4.3 Binary properties
To estimate the masses of the T dwarf companions, we employed the evolutionary tracks of from the Sonora-Bobcat models (Marley et al., 2021), using their estimated ages and the average of T dwarfs with corresponding spectral types. As illustrated in Fig. 10, L 122-88 B has an estimated mass of 0.05–0.067 M for an age of 2.5-9.9 Gyr, while UPM J10403551 Ba and Bb have estimated masses of 0.012–0.04 M and 0.007–0.02 M⊙, respectively, for an age of 0.3-2.0 Gyr. Using the derived distances, projected separations, and masses of all components, we calculated the binding energies of L 122-88 AB and UPM J10403551 AabBab to be and J, respectively.
@@
We found that L 122-88 AB and UPM J10403551 AabBab have values of 1.02 and 0.97 pc, respectively. Their projected separations correspond to 0.034 and 0.008 , respectively, which are far smaller than their respective values. This analysis strongly suggests that both systems are gravitationally stable. The spectroscopic distances of L 122-88 B and UPM J10403551 B are also consistent to Gaia parallax distances of their primaries (see Table 1).